# myboxenhanced,colback=blue!5!white,colframe=blue!75!black,boxrule=0.6mm,fontupper=

### v1

@@ 7 Conclusion and Future Work
Understanding the intricacies of Reinforcement Learning (RL) ’s complexities is challenging, exacerbated by the divergent methodologies employed across different problem domains. This fragmentation hinders the development of unifying principles and consistent practices in RL. To address this critical gap, we propose an innovative framework to understand different methods of effectively integrating the inherent structure of learning problems into RL algorithms. Our work serves as a pivotal step towards consolidating the multifaceted aspects of RL, ushering in a design pattern perspective for this domain.
@@ Acknowledgments
The authors wish to thank Robert Kirk and Rohan Chitnis for their discussion and comments on drafts of this work.

### v3

@@ 8 Conclusion and Future Work
Understanding the intricacies and complexities of Deep RL is challenging, exacerbated by the divergent methodologies employed across different problem domains. This fragmentation hinders the development of unifying principles and consistent practices in RL. To address this critical gap, we propose an innovative framework to understand different methods of effectively integrating the inherent structure of learning problems into RL algorithms. Our work serves as a pivotal step towards consolidating the multifaceted aspects of RL, ushering in a design pattern perspective for this domain.
@@ Acknowledgments
We wish to thank Robert Kirk and Rohan Chitnis for their discussion and comments on drafts of this work. We would also like to thank Vincent François-Lavet, Khimya Khetarpal, and Rishabh Aggarwal for providing additional relevant references in the literature.